# A Quadratic Programming Approach to Flight Envelope Protection Using Control Barrier Functions

### v2

@@ 5.2 Single Operating Point FEP Performance under Disturbances and Parametric Uncertainties
While the previous results focused on the nominal performance of FEP strategies, the following part investigates the robustness of CBF-based safety filters in the presence of external disturbances and model uncertainties. It is important to note that although CBFs are a powerful tool that provides formal guarantees on forward invariance, these guarantees fundamentally depend on the accuracy of the underlying system model. In real-world scenarios, however, discrepancies between the model and the actual system behavior, caused by disturbances, degradation, or unmodeled dynamics, can cause the system to violate safety constraints. This limitation is inherent in all model-based techniques. Adaptive CBF approaches, such as those proposed in [16, 32], aim to restore safety guarantees through online adaptation of model parameters. However, the integration of such adaptive schemes is beyond the scope of this study. Instead, during the analysis of the robustness properties, it was observed that standard CBF safety filters are unable to maintain FEP when the system is subject to external disturbances or significant parametric uncertainties. To address this issue, we utilize the inherent connection between CBFs and Control Lyapunov Functions (CLFs), as proposed by [33]. Specifically, in situations where the system leaves the safe set, the barrier function $h_{i}(x)$ associated with the respective HOCBF constraint can be reinterpreted as a stabilizing objective by defining an auxiliary CLF:
@@
While disturbances primarily account for real-time deviations from the nominal trajectory, parametric uncertainties represent structural mismatches between the modeled and actual system behavior. These may arise from outdated aerodynamic data, actuator modeling errors, or changes in environmental conditions. To model parametric uncertainties, static deviations in key aerodynamic coefficients are considered as: $\Delta C_{Z,\alpha}=-40\%$,$\Delta C_{Z,\delta}=-30\%$, and $\Delta C_{m,q}=-10\%$.
@@ 5.3 Extended FEP performance analysis
[…]
@@
Figure 12 presents the results of this case. As illustrated in the upper left plot of Figure 12, the applied reference command $\alpha_{r}(t)$ (black dash-dotted line) is sinusoidal and intentionally violates the imposed safety constraints, including the stall limit of ±$\pm 15^{\circ}$ and load factor constraints of ±$\pm 10$. Initially, the stall constraints are most critical due to low dynamic pressure, while load factor limitations dominate later phases as the vehicle accelerates. Despite these challenges, the proposed CBF-based safety filter successfully ensures strict satisfaction of the considered state constraints throughout the entire flight envelope. The lower subplots of Figure 12 show the time evolution of additional flight variables. Notably, both barrier values, $h_{1}$ and $h_{2}$, remain strictly positive, confirming the forward invariance of the safe set. The actuator input $\delta(t)$ remains within admissible bounds, illustrating that the safety filter accounts for input constraints while maintaining closed-loop performance.

### v3

@@ 5.2 Single Operating Point FEP Performance under Disturbances and Parametric Uncertainties
While the previous results focused on the nominal performance of FEP strategies, the following part investigates the robustness of CBF-based safety filters in the presence of external disturbances and model uncertainties. It is important to note that although CBFs are a powerful tool that provides formal guarantees on forward invariance, these guarantees fundamentally depend on the accuracy of the underlying system model. In real-world scenarios, however, discrepancies between the model and the actual system behavior, caused by disturbances, degradation, or unmodeled dynamics, can cause the system to violate safety constraints. This limitation is inherent in all model-based techniques. Adaptive CBF approaches, such as those proposed in [Taylor2020-em, Autenrieb2023b], aim to restore safety guarantees through online adaptation of model parameters. However, the integration of such adaptive schemes is beyond the scope of this study. Instead, during the analysis of the robustness properties, it was observed that standard CBF safety filters are unable to maintain FEP when the system is subject to external disturbances or significant parametric uncertainties. To address this issue, we utilize the inherent connection between CBFs and Control Lyapunov Functions (CLFs), as proposed by [Catellani_2024]. Specifically, in situations where the system leaves the safe set, the barrier function $h_{i}(x)$ associated with the respective HOCBF constraint can be reinterpreted as a stabilizing objective by defining an auxiliary CLF:
@@
While disturbances primarily account for real-time deviations from the nominal trajectory, parametric uncertainties represent structural mismatches between the modeled and actual system behavior. These may arise from outdated aerodynamic data, actuator modeling errors, or changes in environmental conditions. To model parametric uncertainties, static deviations in key aerodynamic coefficients are considered as: $\Delta C_{Z,\alpha}=-40\%$$\Delta C_{Z,\delta}=-30\%$, and $\Delta C_{m,q}=-10\%$.
@@ 5.3 Extended FEP performance analysis
To further validate the proposed approach across a wide range of operating conditions, an extended simulation is conducted over a 150-second trajectory covering a flight velocity range from $V=100\,\text{m/s}$ to $V=1500\,\text{m/s}$. In this scenario, the complete flight envelope is explored while both stall and load factor constraints on the angle of attack are considered, as discussed in Section 3.
@@
Figure 12 presents the results of this case. As illustrated in the upper left plot of Figure 12, the applied reference command $\alpha_{r}(t)$ (black dash-dotted line) is sinusoidal and intentionally violates the imposed safety constraints, including the stall limit of $\pm 15^{\circ}$ and load factor constraints of $\pm 10$. Initially, the stall constraints are most critical due to low dynamic pressure, while load factor limitations dominate later phases as the vehicle accelerates. Despite these challenges, the proposed CBF-based safety filter successfully ensures strict satisfaction of the considered state constraints throughout the entire flight envelope. The lower subplots of Figure 12 show the time evolution of additional flight variables. Notably, both barrier values, $h_{1}$ and $h_{2}$, remain strictly positive, confirming the forward invariance of the safe set. The actuator input $\delta(t)$ remains within admissible bounds, illustrating that the safety filter accounts for input constraints while maintaining closed-loop performance.